August 6th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCal
      • still cabling - 9 bundles
      • 32 cables additionnal + 32 spares channels
      • slack to be moved to HCAL side : difficult process on-going - might take a few days
      • hope to be done next week
    • Bigbite Shower
      • finished HV cables
      • working signal cables and trigger monitor
      • preshower finished on BB, shower cabling but others can use platform
    • Bigbite Hodoscope
      • were waiting for cable trays - in place now - will start lying cables ADC and TDC
      • software progress
      • Level translator can be powered
    • GRINCH
      • moved all slack near BigBite
      • HV cables : patch, can install long HV cables , HV cables at top of BB, need lift and will install TDC at same side
    • LHRS
      • almost ready, a few missing channels
      • Bob will help Ole in counting room - 2 weeks - might need more people
    • CDet
      • student working on analysis with new students
      • working on resistors and taking data
      • 1.5 module left to go
      • some PMT ordered
    • GEM BB Hall A
      • gas flowing on BB side
      • LV connection on-going
      • HV ready
      • checking connection : MPD fine, possibly one broken fiber
      • few spare lines avalaible
      • almost ready for readout
      • shielding : not solid plan for INFN in BigBite - could not run HV on chamber - making improvements adding kapton - ~2 days of trouble shooting
      • SSP readout almost done with CODA - should be avalaible at end of day - running on adaq1
      • need to update SSP in Hall A
    • GEM INFN Test Lab
      • J1 all shielding, J3 half installed
      • SSP readout in INFN all up to date and running
    • GEM UVA
    • Beamline
      • working on live plots for shift takers
      • hacsbc2
      • run copy to SBS bunker
    • Helicity
      • helicity scaler installed on LHRS - was working in test setup
    • RF time
      • need to connect in LHRS and send copy to
    • HV
      • rpi online : GRINCH and shower should be ok
      • BB hodoscope need to be powered - same extension
      • MPOD installed - need to set it up on network
      • EPICS software avalaible for all
      • rpi17 and rpi18 off now - but can be turned one
    • Scalers
      • crate for scalers installed
      • working on software to handle different scalers
      • can turn on FADC scalers but depends on thresholds
      • scaler server run during run and between run
      • shower scalers can go in scaler crate if not there already
    • VME controllers
      • all VME controllers booted for shower, hodoscope and GEM, GRINCH



  • On-going tasks
    • Network setup ( done ) - SBS MPD side might need few cables
    • Scaler crate setup (50 % - Steve ) - need fix supply
    • documentation of VME crate names to power cycle
    • 3 rd GEM VXS crate install router moved up ( 50% )
    • need power and network in GEM bunker ( done ) - need connect computer
    • detector accounts - will just use vncservers on different machines -
    • boot VME controllers (90 % done - need ask hostname TS crate )
    • TS cables - all fibers plugged - will post informations -hodoscope / grinch/ shower separate - can be combined at any time now everything plugged to TS
    • detector cabling
    • new MPD production - about to start - FPGA shipped to Paolo
    • MPD VTP readout - on-going in daq lab - still having issue getting APV sending data to MPD - events going through high speed link
    • MPD being setup in Hall A - need setup VTPs to boot - 10 gigE network plugged - will be able to test in Hall A - LV installed soon
    • setup portservers : 1 4 port portserver for VTP, still need to set up a 8 port one and HCAL portserver
    • need to install remote switch for crate in bunker - rest are VXS or VME64 -
    • Bigbite - HCal trigger setup - need check with Bob if additionnal NIM crate need to be install
    • TS setup - 50 %
    • Beamline crate setup ( in scaler crate ) - FADC is in scaler need to be setup
    • HCAL cabling
    • HCAL need power ( done, Berta )
    • Mark hodoscope trigger ?
    • Computer move to Hall A : sbsdaq1/2 tedbbdaq (done ) enpcamsonne ( will move next week )
    • LHRS fiber pull ( waiting fiber August 13th ) - most likely LHRS at end of SBS run
    • BPM/BCM cables SBS weldment ( will do next week )
  • Outstanding issues
    • 2 bad FADCs Bigbite - 1 FADC was repaired - one being repaired - 1 Compton FADC - repaired can use as spare ( done )
    • SD delay : still need 2 SDs ( Glasgow and INFN just requested quote )
    • need to adjust for HCAL transit time with different PMTs
    • status of cosmics paddles ( installed by Bogdan )
    • GEM fibers too short : Mark will order parts
    • 15 bad MPDs - need to look at fixing - Ben ? Dustin loaning us 6 MPDs - need 85 total - 67 in hand
      • APV needed 1208 - 966 in hand working - Michael might lend 40 - ( check with Thir )
    • GEM fiber possibly bad
    • GEM HV IP need setup ( Alex )